# APCTP-Pre2023-010

### v2

@@ V.2 A toy model with massless isocurvature perturbations
$\mathcal{M}_{\rm bb}=T^{2}\quad\text{and}\quad\mathcal{M}_{\rm nb}=(3-\epsilon)T\,,$ (137)
thus fulfilling the first condition of (136). Demanding the vanishing of the second condition fixes the last component of the mass matrix to $\mu_{\rm eff}=0$ or $m_{\rm eff}=-3T^{2}$, akin to the two-field ultra-light scenario Achucarro:2016fby where the torsion is identically zero and the effective mass of the orthogonal field is also zero, $m_{\rm eff}=0$. Having fixed the geometric properties of the field space, we now turn our attention to the potential.
To understand the conditions that can lead to such a potential we will use the effective potential construction Christodoulidis:2019mkj . This assumes that a number of fields follow their potential gradient while the rest fields are frozen at the minima of their effective potential. As shown in Christodoulidis:2022vww , a non-zero torsion in three-field models is possible when considering e.g. a diagonal metric
$G_{IJ}=\text{diag}(f^{2},g^{2},1)\,,$ (138)
@@
$V_{,\chi\chi}+3{f_{,\chi}^{2}\over f^{4}}(w_{1})^{2}H^{2}+3{g_{,\chi}^{2}\over g^{4}}(w_{2})^{2}H^{2}-{f_{,\chi\chi}\over f^{3}}(w_{1})^{2}H^{2}-{g_{,\chi\chi}\over g^{3}}(w_{2})^{2}H^{2}=0\,,$ (140)
@@
${f_{,\chi}^{2}\over f^{4}}(w_{1})^{2}+{g_{,\chi}^{2}\over g^{4}}(w_{2})^{2}$ $\approx T^{2}+\Omega^{2}\,,$ (141)
${f_{,\chi\chi}\over f^{3}}(w_{1})^{2}H^{2}+{g_{,\chi\chi}\over g^{3}}(w_{2})^{2}H^{2}$ $\approx-R_{{\rm n}IJ{\rm n}}\dot{\phi}^{I}\dot{\phi}^{J}\,,$ (142)
we find that (140) is exactly equal to $\mu_{\rm eff}=0$. The problem has now reduced to finding a potential that satisfies (139) and (140).
The simplest way is to consider that (139) is satisfied identically, integrating which implies that the potential satisfies
$V+{1\over 2}{(w_{1})^{2}H^{2}\over f^{2}}+{1\over 2}{(w_{2})^{2}H^{2}\over g^{2}}=c(\phi^{1},\phi^{2})\,.$ (143)
@@
$w_{I}=2{H_{,I}\over H}-{H^{2}G^{KJ}w_{KI}w_{J}\over 3-\epsilon}\,,$ (144)
and discarding the last term [see the discussion around (130)] we finally obtain the form of the potential
$V=3H^{2}-2G^{IJ}H_{,I}H_{,J}\,,$ (145)
with $H$ being independent of $\chi$. Any such potential will lead to massless isocurvature perturbations on the attractor solution. More generally, the previous derivation is an example of the superpotential method in inflation (see also Achucarro:2019pux ; Aragam:2019omo for multi-field constructions). In this method it is assumed that the velocities in the evolution of the Hubble parameter are given as
$\dot{\phi}^{I}=-2G^{IJ}{H_{,J}}\,,$ (146)
and the form of the potential follows after substituting this choice into the Friedman constraint. It can be checked by direct substitution that the mass matrix elements for this specific expression of the potential agree with the expressions (137) and $\mu_{\rm eff}=0$. Finally, making a choice for the field metric and the Hubble parameter $H(\phi^{1},\phi^{2})$ leads to specific values of the background quantities $\epsilon$, $\Omega$ and $T$.
@@
$P_{\mathcal{R}}=\mathcal{R}^{\prime}-\sqrt{\frac{2}{\epsilon}}\Omega q_{\rm n}=0\,.$ (147)
@@
$q_{\rm n}^{\prime}=0\,,\qquad P_{\rm b}^{\prime}=0\,,\qquad P_{\rm b}=-Tq_{\rm n}\,,\qquad P_{\rm n}=-Tq_{\rm b}\,,$ (148)
with $q_{\rm n}$ and $q_{\rm b}$ being non-zero, implying that the two orthogonal perturbations are unstable. Note that this does not contradict the superhorizon analysis of section IV.2 where it was found that the torsion in general has a stabilizing effect because here multiple eigenvalues have zero real part and any arguments based on linearization fail. Regarding the curvature perturbation the relation (147) still holds in the presence of torsion and we obtain the additional relation
$q_{\rm b}^{\prime}=-2Tq_{\rm n}\,.$ (149)
@@
In figure 3 we verify numerically our previous arguments regarding the behaviour of the three perturbations and their corresponding momenta (which we do not plot here) for a model with $\epsilon=0.01$ and for two sets of the bending parameters: $\Omega=\sqrt{5}\Theta(N-3)$, $T=0$ (dashed) and $\Omega=2\Theta(N-3)$, $T=\Theta(N-3)$ (solid), where $\Theta(\cdots)$ is the step function. Notice also that for both sets of parameters the exponential amplification on subhorizon scales begins when the inequality (135) is satisfied. We have chosen the turn rate and torsion such that $\Omega^{2}+T^{2}$ is the same for both models and we observe that for the two models the curvature and normal perturbations behave almost identically (see also the discussion at the end of section V.1). However, the second model predicts sizable entropic perturbations at the end of inflation.
Figure 3: The power spectra (normalized by the single-field spectrum ${\cal{P}}_{0}$) for a given mode that crosses the horizon at $N=0$. The turning rate and torsion are zero before $N=-3$, and become constant afterwards. The three perturbations, $\mathcal{R}$, $q_{\rm n}/\sqrt{2\epsilon}$ and $q_{\rm b}/\sqrt{2\epsilon}$, all with the Bunch-Davies initial conditions, are depicted in thin dashed purple, red and gray respectively for the model without torsion, and in thick solid blue, orange and black respectively for the model with torsion. After the horizon crossing the normal perturbation $q_{\rm n}$ freezes whereas the curvature perturbation $\mathcal{R}$ continues to increase. The behaviour of the binormal perturbation $q_{\rm b}$ depends on the value of the torsion: For the zero torsion case it is effectively decoupled from the other two and behaves as a single-field perturbation, whereas for the non-zero case it continues to increase on superhorizon scales. The vertical line at $N\approx-1.5$ corresponds to the $e$-folding time when (135) is satisfied and the exponential amplification of perturbations commences.
VI Conclusions
@@
More specifically, in section II we introduced the basic ingredients of multi-field perturbations and showed that a simple change of the variable $x\equiv-k\tau$ and a field redefinition by $\tilde{v}_{i}=\sqrt{k}v_{i}$ trivially yields the scale dependence of the power spectrum when every background quantity (except for $\epsilon$ and $\eta$) is constant. The power spectrum for these models becomes identical to that of single-field models, a result which follows simply by inspecting the equations for perturbations without resorting to an effective field theory or to the slow-turn limit. We also pointed out that the behaviour of perturbations becomes more transparent in the Hamiltonian description. Then, in section III we quantified the enhancement of perturbations for two-field systems that undergo constant turns of short and long duration, explaining how the two cases have to be treated differently. More specifically, for sharp turns the solution is found by exponentiating the stability matrix of perturbations, while for broad turns one has to first write the system into perturbative expansion over $1/\Omega$ and then try to solve it perturbatively. In the latter case for solutions of the form $A\exp(B)$ the phase $B$, which can be either imaginary or real, is found by solving the part of the equations without the Hubble friction terms, while the correct amplitude is found only once the Hubble friction terms are taken into account. Although we have limited our discussion on two fields, it is straightforward to generalize our results for more fields.
@@
Subsequently, in section IV we investigated under what general conditions the curvature perturbation becomes exponentially enhanced on subhorizon scales, and when the perturbations remain stable on superhorizon scales, listing a series of conditions. We applied our findings on three-field models with isometries in section V and showed that the amplification happens when the quantity $\mathcal{M}_{\rm nn}-\Omega^{2}-T^{2}$ becomes negative, very analogous to the two-field case. We continued with a discussion of the ultra-light scenario with all three fields being massless on superhorizon scales and showed that when the torsion of the trajectory in the field space is non-zero then, in addition to the curvature perturbation, the binormal perturbation is also continuously sourced by the normal perturbation, the latter being constant. This partly answers the question of whether the presence of the torsion can produce distinct observational signatures in models with more than two fields. We also refer the interested readers to Aragam:2023adu where the authors studied the amplification of the power spectrum for three-field systems due to a sudden increase of the turn rate and torsion.
Acknowledgements.
@@
It is a pleasure to thank George Kodaxis, Maria Mylova and Nikolaos Tetradis for useful discussions at the early stages of this work. We would also like to thank Gonzalo Palma and Spyros Sypsas for useful discussions during the International Molecule-Type Workshop “Revisiting cosmological non-linearities in the era of precision surveys (YITP-T-23-03)” at the Yukawa Institute for Theoretical Physics, Kyoto, Japan. We are supported in part by the National Research Foundation of Korea Grant 2019R1A2C2085023. JG also acknowledges the Korea-Japan Basic Scientific Cooperation Program supported by the National Research Foundation of Korea and the Japan Society for the Promotion of Science (2020K2A9A2A08000097). JG is further supported in part by the Ewha Womans University Research Grant of 2022 (1-2022-0606-001-1) and 2023 (1-2023-0748-001-1). JG is also grateful to the Asia Pacific Center for Theoretical Physics for hospitality while this work was under progress.
@@ Appendix A Massive field in quasi-de Sitter space
The curvature perturbation in single-field inflation is a massless scalar field in the sense that it freezes on superhorizon scales. To find the approximate analytic solution it is better to switch to the canonical field $u=\sqrt{2\epsilon}\mathcal{R}$ which obeys the Mukhanov-Sasaki equation Sasaki:1986hm ; Mukhanov:1988jd :
$u_{,\tau\tau}+\left[k^{2}-(aH)^{2}\left(2-\epsilon+{3\over 2}\eta-{1\over 2}\epsilon\eta+{1\over 4}\eta^{2}+{1\over 2}\eta\xi\right)\right]u=0\,,$ (150)
@@
$\tau-\tau_{\rm r}=\int^{N}_{N_{\rm r}}\,\mathrm{d}{N}^{\prime}{1\over aH}=\int^{N}_{N_{\rm r}}\,\mathrm{d}{N}^{\prime}{1\over a_{0}}e^{-{N}^{\prime}+\int\,\mathrm{d}n\epsilon(n)}\,,$ (151)
@@
$\tau-\tau_{\rm r}=-{1+\epsilon\over aH}\Big{|}^{N}_{N_{\rm r}}+\int\,\mathrm{d}{N}^{\prime}{\epsilon^{2}+\epsilon\eta\over aH}\,,$ (152)
where the second term is at least one order higher in the slow-roll parameters. [^10] By successive integration by parts one can derive the relation of the conformal time in terms of the Hubble radius to arbitrary order in the slow-roll parameters. To first order in the slow-roll parameters (150) becomes
$u_{,\tau\tau}+\left[k^{2}-{1\over\tau^{2}}\left(2+3\epsilon+{3\over 2}\eta\right)\right]u=0\,,$ (153)
with the solutions being the Hänkel functions of first and second type. Choosing the positive frequency solution at early times yields
$u={1\over 2}\sqrt{\pi}e^{i(2\nu+1)\pi/4}\sqrt{-\tau}H^{(1)}_{\nu}(\tau)\,,$ (154)
@@
$\nu=\sqrt{{9\over 4}-\mathcal{M}_{\rm nn}}\,.$ (155)

### v3

@@ 5.2 A toy model with massless isocurvature perturbations
$\mathcal{M}_{\rm bb}=T^{2}\quad\text{and}\quad\mathcal{M}_{\rm nb}=(3-\epsilon)T\,,$ (5.11)
thus fulfilling the first condition of (5.10). Demanding the vanishing of the second condition fixes the last component of the mass matrix to $\mu_{\rm eff}=0$ or $m_{\rm eff}=-3T^{2}$, akin to the two-field ultra-light scenario [73] where the torsion is identically zero and the effective mass of the orthogonal field is also zero, $m_{\rm eff}=0$. Having fixed the geometric properties of the field space, we now turn our attention to the potential.
To understand the conditions that can lead to such a potential we will use the effective potential construction [74]. This assumes that a number of fields follow their potential gradient while the rest fields are frozen at the minima of their effective potential. As shown in [72], a non-zero torsion in three-field models is possible when considering e.g. a diagonal metric
$G_{IJ}=\text{diag}(f^{2},g^{2},1)\,,$ (5.12)
@@
$V_{,\chi\chi}+3{f_{,\chi}^{2}\over f^{4}}(w_{1})^{2}H^{2}+3{g_{,\chi}^{2}\over g^{4}}(w_{2})^{2}H^{2}-{f_{,\chi\chi}\over f^{3}}(w_{1})^{2}H^{2}-{g_{,\chi\chi}\over g^{3}}(w_{2})^{2}H^{2}=0\,,$ (5.14)
@@
${f_{,\chi}^{2}\over f^{4}}(w_{1})^{2}+{g_{,\chi}^{2}\over g^{4}}(w_{2})^{2}$ $\approx T^{2}+\Omega^{2}\,,$ (5.15)
${f_{,\chi\chi}\over f^{3}}(w_{1})^{2}H^{2}+{g_{,\chi\chi}\over g^{3}}(w_{2})^{2}H^{2}$ $\approx-R_{{\rm n}IJ{\rm n}}\dot{\phi}^{I}\dot{\phi}^{J}\,,$ (5.16)
we find that (5.14) is exactly equal to $\mu_{\rm eff}=0$. The problem has now reduced to finding a potential that satisfies (5.13) and (5.14).
The simplest way is to consider that (5.13) is satisfied identically, integrating which implies that the potential satisfies
$V+{1\over 2}{(w_{1})^{2}H^{2}\over f^{2}}+{1\over 2}{(w_{2})^{2}H^{2}\over g^{2}}=c(\phi^{1},\phi^{2})\,.$ (5.17)
@@
$w_{I}=2{H_{,I}\over H}-{H^{2}G^{KJ}w_{KI}w_{J}\over 3-\epsilon}\,,$ (5.18)
and discarding the last term [see the discussion around (5.4)] we finally obtain the form of the potential
$V=3H^{2}-2G^{IJ}H_{,I}H_{,J}\,,$ (5.19)
with $H$ being independent of $\chi$. Any such potential will lead to massless isocurvature perturbations on the attractor solution. More generally, the previous derivation is an example of the superpotential method in inflation (see also [75, 76] for multi-field constructions). In this method it is assumed that the velocities in the evolution of the Hubble parameter are given as
$\dot{\phi}^{I}=-2G^{IJ}{H_{,J}}\,,$ (5.20)
and the form of the potential follows after substituting this choice into the Friedman constraint. It can be checked by direct substitution that the mass matrix elements for this specific expression of the potential agree with the expressions (5.11) and $\mu_{\rm eff}=0$. Finally, making a choice for the field metric and the Hubble parameter $H(\phi^{1},\phi^{2})$ leads to specific values of the background quantities $\epsilon$, $\Omega$ and $T$.
@@
$P_{\mathcal{R}}=\mathcal{R}^{\prime}-\sqrt{\frac{2}{\epsilon}}\Omega q_{\rm n}=0\,.$ (5.21)
@@
$q_{\rm n}^{\prime}=0\,,\qquad P_{\rm b}^{\prime}=0\,,\qquad P_{\rm b}=-Tq_{\rm n}\,,\qquad P_{\rm n}=-Tq_{\rm b}\,,$ (5.22)
with $q_{\rm n}$ and $q_{\rm b}$ being non-zero, implying that the two orthogonal perturbations are unstable. Note that this does not contradict the superhorizon analysis of section 4.2 where it was found that the torsion in general has a stabilizing effect because here multiple eigenvalues have zero real part and any arguments based on linearization fail. Regarding the curvature perturbation the relation (5.21) still holds in the presence of torsion and we obtain the additional relation
$q_{\rm b}^{\prime}=-2Tq_{\rm n}\,.$ (5.23)
@@
In figure 3 we verify numerically our previous arguments regarding the behaviour of the three perturbations and their corresponding momenta (which we do not plot here) for a model with $\epsilon=0.01$ and for two sets of the bending parameters: $\Omega=\sqrt{5}\Theta(N-3)$, $T=0$ (dashed) and $\Omega=2\Theta(N-3)$, $T=\Theta(N-3)$ (solid), where $\Theta(\cdots)$ is the step function. Notice also that for both sets of parameters the exponential amplification on subhorizon scales begins when the inequality (5.9) is satisfied. We have chosen the turn rate and torsion such that $\Omega^{2}+T^{2}$ is the same for both models and we observe that for the two models the curvature and normal perturbations behave almost identically (see also the discussion at the end of section 5.1). However, the second model predicts sizable entropic perturbations at the end of inflation.
Figure 3: The power spectra (normalized by the single-field spectrum ${\cal{P}}_{0}$) for a given mode that crosses the horizon at $N=0$. The turning rate and torsion are zero before $N=-3$, and become constant afterwards. The three perturbations, $\mathcal{R}$, $q_{\rm n}/\sqrt{2\epsilon}$ and $q_{\rm b}/\sqrt{2\epsilon}$, all with the Bunch-Davies initial conditions, are depicted in thin dashed purple, red and gray respectively for the model without torsion, and in thick solid blue, orange and black respectively for the model with torsion. After the horizon crossing the normal perturbation $q_{\rm n}$ freezes whereas the curvature perturbation $\mathcal{R}$ continues to increase. The behaviour of the binormal perturbation $q_{\rm b}$ depends on the value of the torsion: For the zero torsion case it is effectively decoupled from the other two and behaves as a single-field perturbation, whereas for the non-zero case it continues to increase on superhorizon scales. The vertical line at $N\approx-1.5$ corresponds to the $e$-folding time when (5.9) is satisfied and the exponential amplification of perturbations commences.
6 Conclusions
@@
More specifically, in section 2 we introduced the basic ingredients of multi-field perturbations and showed that a simple change of the variable $x\equiv-k\tau$ and a field redefinition by $\tilde{v}_{i}=\sqrt{k}v_{i}$ trivially yields the scale dependence of the power spectrum when every background quantity (except for $\epsilon$ and $\eta$) is constant. The power spectrum for these models becomes identical to that of single-field models, a result which follows simply by inspecting the equations for perturbations without resorting to an effective field theory or to the slow-turn limit. We also pointed out that the behaviour of perturbations becomes more transparent in the Hamiltonian description. Then, in section 3 we quantified the enhancement of perturbations for two-field systems that undergo constant turns of short and long duration, explaining how the two cases have to be treated differently. More specifically, for sharp turns the solution is found by exponentiating the stability matrix of perturbations, while for broad turns one has to first write the system into perturbative expansion over $1/\Omega$ and then try to solve it perturbatively. In the latter case for solutions of the form $A\exp(B)$ the phase $B$, which can be either imaginary or real, is found by solving the part of the equations without the Hubble friction terms, while the correct amplitude is found only once the Hubble friction terms are taken into account. Although we have limited our discussion on two fields, it is straightforward to generalize our results for more fields.
@@
Subsequently, in section 4 we investigated under what general conditions the curvature perturbation becomes exponentially enhanced on subhorizon scales, and when the perturbations remain stable on superhorizon scales, listing a series of conditions. We applied our findings on three-field models with isometries in section 5 and showed that the amplification happens when the quantity $\mathcal{M}_{\rm nn}-\Omega^{2}-T^{2}$ becomes negative, very analogous to the two-field case. We continued with a discussion of the ultra-light scenario with all three fields being massless on superhorizon scales and showed that when the torsion of the trajectory in the field space is non-zero then, in addition to the curvature perturbation, the binormal perturbation is also continuously sourced by the normal perturbation, the latter being constant. This partly answers the question of whether the presence of the torsion can produce distinct observational signatures in models with more than two fields. We also refer the interested readers to [45] where the authors studied the amplification of the power spectrum for three-field systems due to a sudden increase of the turn rate and torsion.
Acknowledgments
@@
It is a pleasure to thank George Kodaxis, Maria Mylova and Nikolaos Tetradis for useful discussions at the early stages of this work. We would also like to thank Gonzalo Palma and Spyros Sypsas for useful discussions during the International Molecule-Type Workshop “Revisiting cosmological non-linearities in the era of precision surveys (YITP-T-23-03)” at the Yukawa Institute for Theoretical Physics, Kyoto, Japan. We are supported in part by the National Research Foundation of Korea Grant 2019R1A2C2085023. JG also acknowledges the Korea-Japan Basic Scientific Cooperation Program supported by the National Research Foundation of Korea and the Japan Society for the Promotion of Science (2020K2A9A2A08000097). JG is further supported in part by the Ewha Womans University Research Grant of 2022 (1-2022-0606-001-1) and 2023 (1-2023-0748-001-1). JG is also grateful to the Asia Pacific Center for Theoretical Physics for its hospitality while this work was in progress.
@@ Appendix A Massive field in quasi-de Sitter space
The curvature perturbation in single-field inflation is a massless scalar field in the sense that it freezes on superhorizon scales. To find the approximate analytic solution it is better to switch to the canonical field $u=\sqrt{2\epsilon}\mathcal{R}$ which obeys the Mukhanov-Sasaki equation [77, 78]:
$u_{,\tau\tau}+\left[k^{2}-(aH)^{2}\left(2-\epsilon+{3\over 2}\eta-{1\over 2}\epsilon\eta+{1\over 4}\eta^{2}+{1\over 2}\eta\xi\right)\right]u=0\,,$ (A.1)
@@
$\tau-\tau_{\rm r}=\int^{N}_{N_{\rm r}}\,\mathrm{d}{N}^{\prime}{1\over aH}=\int^{N}_{N_{\rm r}}\,\mathrm{d}{N}^{\prime}{1\over a_{0}}e^{-{N}^{\prime}+\int\,\mathrm{d}n\epsilon(n)}\,,$ (A.2)
@@
$\tau-\tau_{\rm r}=-{1+\epsilon\over aH}\Big{|}^{N}_{N_{\rm r}}+\int\,\mathrm{d}{N}^{\prime}{\epsilon^{2}+\epsilon\eta\over aH}\,,$ (A.3)
where the second term is at least one order higher in the slow-roll parameters. [^10] By successive integration by parts one can derive the relation of the conformal time in terms of the Hubble radius to arbitrary order in the slow-roll parameters. To first order in the slow-roll parameters (A.1) becomes
$u_{,\tau\tau}+\left[k^{2}-{1\over\tau^{2}}\left(2+3\epsilon+{3\over 2}\eta\right)\right]u=0\,,$ (A.4)
with solutions the Hänkel functions of first and second type. Choosing the positive frequency solution at early times yields
$u={1\over 2}\sqrt{\pi}e^{i(2\nu+1)\pi/4}\sqrt{-\tau}H^{(1)}_{\nu}(\tau)\,,$ (A.5)
@@
$\nu=\sqrt{{9\over 4}-\mathcal{M}_{\rm nn}}\,.$ (A.6)
Appendix B Sub- to superhorizon evolution for models with $m_{\rm eff}\geq 4\Omega^{2}\gg 1$
In this section we will examine how our previous method works for models when a moderate amplification on subhorizon scales is possible.
B.1 $m_{\rm eff}>4\Omega^{2}$
First, we mention models where the mass parameter is strictly larger than $4\Omega^{2}$, which is equivalent to a positive effective mass on subhorizon scales $m_{\rm sub}^{2}=m_{\rm eff}-4\Omega^{2}>0$. At scales $x\sim\sqrt{m_{\rm eff}}$ the two perturbations start deviating from the Bunch-Davies values (assuming they are initially in such states) and decrease at a faster rate than the Bunch-Davies solution. The curvature perturbation freezes on superhorizon scales but approaches its asymptotic value faster than single-field models, behaving as a single-field perturbation in the limit $m_{\rm eff}\rightarrow\infty$, see e.g. [49, 51, 61, 56]. Since these models do not yield any significant amplification, we ignore them and proceed to the borderline case.
B.2 The borderline case $m_{\rm eff}=4\Omega^{2}$
Following the same steps as in Section 3.2 we notice a peculiarity for the borderline mass case $m_{\rm eff}=4\Omega^{2}$. Splitting the stability matrix as in (3.46), i.e. writing
$Y_{,\theta}=(J_{\rm D}+P_{\rm D})Y+P_{\rm od}Y\,,\,$ (B.1)
we find that the off-diagonal terms for the rotated system, represented by the matrix $P_{\rm od}$, become $\mathcal{O}(1)$ when $x\sim\sqrt{2\Omega}$ and then keep increasing until horizon crossing. Thus, we can not trust this approximation after this scale. Note also that choosing to diagonalize the full stability matrix instead of its zeroth-order part does not alter the previous behaviour. Using the formulae listed in Section 3.2 we find that at $x\sim\sqrt{2\Omega}$ the four variables have almost equal amplitudes $\sim\Omega^{1/4}$ modulo some $\mathcal{O}(1)$ constants. Since on scales $x\sim\sqrt{2\Omega}$ the quadratic terms $x^{2}/(2\Omega)^{2}$ in (3.42)-(3.45) become $\mathcal{O}(\Omega^{-1})$, this implies they can be moved to the stability matrix of the first-order part, suggesting the following split:
$\tilde{J}_{(0)}\equiv\begin{pmatrix}0&1&1&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&-1&-1&0\end{pmatrix}\quad\text{and}\quad\tilde{J}_{(1)}\equiv\begin{pmatrix}0&0&0&0\\ -\left(\dfrac{x}{2\Omega}\right)^{2}&-\dfrac{3}{2\Omega}&0&0\\ 0&0&0&0\\ 0&0&-\left(\dfrac{x}{2\Omega}\right)^{2}&-\dfrac{3}{2\Omega}\end{pmatrix}\,.$ (B.2)
By doing so we are able to solve the system consistently. In particular, we find that the curvature perturbation freezes at $x\sim\sqrt{2\Omega}$, $|\mathcal{R}|_{x<\sqrt{2\Omega}}\sim|\tilde{p}_{\mathcal{R}}|_{x=\sqrt{2\Omega}}\sim(2\Omega)^{1/4}$ (recovering the result of [56]), whereas the isocurvature perturbation decays as $x^{3/2}/\sqrt{2\Omega}$ (see figure 4).
Figure 4: The curvature (blue) and isocurvature (orange) power spectra normalized by their asymptotic expressions for the borderline case $m_{\rm eff}=4\Omega^{2}$ in the regime $x<\sqrt{2\Omega}$. Each curve is the numerical evolution for different values of the turn rate: solid, dashed and dotted curves represent $\Omega=$20, 50 and 100 respectively. The horizon crossing point is set at $N=0$.